# DeepVoid: A Deep Learning Void Detector

### v1

@@ 5.4 Prediction
DeepVoid computes class scores for void, wall, filament, and halo classes that sum to unity in each voxel and assigns the class with the highest score as the prediction for each voxel (see Section 4.1). Due to GPU memory constraints, we cannot load and perform a prediction on an entire TNG mass density cube. Instead, we split the volume into overlapping subcubes of size $128^{3}$ that overlap by half, similar to the training process. DeepVoid then predicts in batches of eight subcubes. To reconstruct the full predicted volume, we reassemble the overlapping subcubes by selecting only the central region of each predicted subcube. This recombination process produces a seamless, non-overlapping reconstruction of the predicted semantic segmentation with consistent predictions across subcube boundaries.
6 Void-Finding Analysis
@@
To quantitatively evaluate the performance of the DeepVoid models, we employ a combination of standard multi-label classification metrics. All of these metrics are calculated from the confusion matrix (as shown below in Figure 7). The confusion matrix is a fundamental tool for evaluating classifiers that compares predicted class labels against the physics-based truth labels, so diagonal elements in the matrix represent correct classifications, while the non-diagonal elements represent the instances where the predicted and true class do not match. Definitions and a more detailed discussion of these metrics can be found in Appendix B.
To evaluate model performance, we focus on the void class F1 score and the Matthews correlation coefficient (MCC) as the primary indicators of model performance. We also consider balanced accuracy and the micro-averaged F1 score. Balanced accuracy and F1 scores range from zero to one, while the MCC ranges from $-1$ to 1, where a zero MCC corresponds to random guessing. These metrics were selected because of their interpretability and robustness to class imbalance.
6.1 DeepVoid Base Models
As discussed above, we train base models on the full-density TNG-300-3-Dark simulation, with interparticle separation of $\lambda=0.33$ $h^{-1}$ Mpc. The best base TNG model achieves a void F1 score of 0.96 and a MCC of 0.81. Table 1 shows performance metrics of the best DeepVoid base models trained on the TNG DM particles and using SCCE loss functions. The upper row of Figure 5 compares the prediction of the best base model with the truth table from the tidal tensor classification; there is excellent visual agreement. In Section 7.3 below we quantify details of this visual comparison in the context of examining the effects of sparse sampling on the model performance.
@@ 6.1 DeepVoid Base Models
6.2 Alternative Validation Test
As an additional test, we rotate the DM particles/halo distribution in TNG300 by 45 and recompute the tidal tensor mask. CNNs are by nature not rotationally invariant (Zeiler & Fergus, 2013).
@@ 6.2 Alternative Validation Test
7 Training on Sparser Samples
@@
Figure 5: Upper row: TNG mass density contrast and prediction for the best base TNG model predicting on the full TNG particle distribution with $\lambda=0.33$ $h^{-1}$ Mpc. Upper right panel is the tidal tensor truth table, which is the same for all models. Lower row: TNG subhalo positions and prediction for the best CL model. Subhalos are first ordered by mass, then abundance matched such that the average intertracer separation is $\lambda=10$ $h^{-1}$ Mpc. Subhalos that are within 2 $h^{-1}$ Mpc of either side of the slice taken through the simulation are shown. The base model was trained on the TNG full density with a U-net with 3 layers of depth and 32 initial filters. The model was then curricular-trained on subhalos with $\lambda=3$ $h^{-1}$ Mpc average spacing with the first encoding level frozen, then the second encoding level was frozen, and the model was curricular trained out to $\lambda=10$ $h^{-1}$ Mpc.
@@
To explore building a version of DeepVoid that can be applied to density fields from deep galaxy surveys, we must train and predict on density fields with larger intertracer spacings than the full DM particle fields of TNG. The subhalo catalog of TNG provides an adequate test case for DeepVoid, as described in Section 2.
7.1 Directly Training with Sparse Samples
@@
The most naive way to predict the morphology of a volume with large average intertracer separation would be to simply run a prediction using a base model. As expected, the base models, which are trained on density fields with intertracer spacings of much less than one, perform very poorly when predicting on the $\lambda=10$ $h^{-1}$ Mpc subhalo mass density. The best TNG base model reaches a balanced accuracy of 0.27 and an MCC of 0.07. Recalling the definition of the MCC, that score is only slightly better than random guessing.
@@
Next, we train a new model directly on the volumes with intertracer separations similar to that of galaxy redshift surveys of interest. This vastly outperforms the base models. Training on the subhalo density field with $\lambda=10$ $h^{-1}$ Mpc results in scores of void F1=0.89 and MCC=0.56. As discussed in the next section, we can further improve the models’ predictive performance using a multi-step training process.
@@
Figure 6: Slices through the TNG300-3-Dark volume. The tidal tensor mask is shown in gray-scale on both panels. Red voxels represent true void voxels misclassified as wall, while blue voxels show true wall voxels misclassified as void. Left: base model trained on TNG subhalos with a 0.33 $h^{-1}$ Mpc separation. Right: CL model trained on subhalos with a 10 $h^{-1}$ Mpc spacing. As the intertracer separation increases, the lower signal-to-noise ratio negatively impacts the quality of the segmentation, causing the model to miss small-scale features such as walls between smaller voids and tendrils of filaments extending into voids.
@@
Figure 7: Comparison of two confusion matrices generated by running model predictions on the TNG validation dataset. Left: confusion matrix from a model trained with depth of 3 and 32 initial filters on the full DM density (average interparticle spacing of 0.33 $h^{-1}$ Mpc). Right: confusion matrix from a model with depth of 3 and 32 initial filters predicting on TNG subhalos with an average intertracer spacing of 10 $h^{-1}$ Mpc. This model was originally trained on the full TNG DM density, then curricular-trained out to 3 $h^{-1}$ Mpc, and again to 10 $h^{-1}$ Mpc. The numbers in the matrices are normalized by the true instances, such that the rows of each matrix sum to one. The base model achieves a void F1 score of 0.96 and a MCC of 0.81 while the CL model achieves a void F1 score of 0.89 and a MCC of 0.6.
7.2 Applying Curricular Learning
@@
Note. — DeepVoid models that used curriculum learning to extend out to an intertracer separation of 10 $h^{-1}$ Mpc. As always, all listed scores were generated from predicting on the validation data. The type of curricular learning employed in each training is denoted by the column CL Type: ENC_EO meaning every other convolution on the entire encoding side is frozen, ENC_D1 and ENC_D2 representing all convolutions on the encoder side being frozen down to a depth of two and three respectively. A CL type of ENC_D1_ENC_D2 means that the model has been curricular trained twice, once to an intermediate intertracer spacing (here $\lambda_{int}=$ 3 $h^{-1}$ Mpc) then again to 10 $h^{-1}$ Mpc. $\lambda_{\rm base}$ represents the interparticle or intertracer spacing of the density field that the base model was trained on, while $\lambda_{\rm int}$ is the intermediate intertracer spacing, if there is one for each respective curricular learning.
As outlined in Section 5.3, curriculum learning aims to successively train a model on increasingly difficult examples. Typically, when training a network, the weights of the model are initialized randomly and then refined iteratively through the process of stochastic gradient descent. However, this is not the case for curricular learning, as some of the initial weights of the model have been preserved from previously trained models that were trained on a simpler example of the task. The idea is that the model inherits valuable knowledge from earlier stages of training, then further iterates to an even better region of the loss landscape. In practice, this means that when we introduce a more difficult example to the model, such as a sparser mass density field, we freeze some of the weights that were learned by training on the ‘easier’ (in this case, denser) example.
We experiment with different schemes of how to freeze the weights. Commonly, the weights of the encoding side of the U-net are frozen to allow for a stable feature extraction process, while the decoding side is allowed to focus on interpreting the already learned features and applying them to the new more challenging data (Ardalan & Subbian, 2022). We perform curricular learning using three different schemes: freezing weights in the entire encoding side (ENC), freezing weights in only one of the two convolutional blocks on the encoding side (ENC_EO), or freezing weights only down to some layer of depth on the encoding side (ENC_D1/D2). For each CL scheme, we begin with weights from the best base model, freeze the specified set of weights, and train the unfrozen weights using training data with an intertracer separation that is larger than that on which the base model was trained. No models with the ENC scheme outperformed the ENC_EO models, so they are not discussed further in this analysis.
@@
For every model in this table, $\lambda_{\rm pred}=10$ $h^{-1}$ Mpc. The best CL models achieve void F1 score of 0.89 and MCC of 0.60. The two-step model slightly outperforms one-step CL models in balanced accuracy. The lower row of Figure 5 shows the result of training a model using this two-step CL scheme out to intertracer separation of 10 $h^{-1}$ Mpc. As with the base model, there is excellent visual agreement between the predicted class labels and the ground truth derived from the tidal tensor. In Section 7.3 we quantify details of this comparison.
Note that Tables 1 and 2 present only the most successful training schemes. We experimented with numerous other combinations of hyperparameters and loss functions not shown here. These results represent a successful proof-of-concept test of classifying LSS from a sparse distribution, but are by no means exhaustive.
From these results, we conclude that performing CL improves the quality of the segmentation for sparser samples. Using CL techniques, we improve training on subhalo distributions with intertracer spacings as large as $\lambda_{\rm pred}=10$ $h^{-1}$ Mpc. We find that the two-step CL scheme (ENC_D1_ENC_D2) yields slightly better results than a single-step CL scheme. Further exploration of model hyperparameters might yield further improvements.
7.3 Segmentation Performance in Sparse Samples
@@
An increase in the number of misclassified voxels with a reduction in tracer density is highlighted in the slice plots of Figure 6. Voxel misclassification is quantified in the confusion matrix in Figure 7, where the left panel describes results for the base model trained on the full DM distribution in TNG300 and the right panel describes results for CL training on very sparse subhalos. The off-diagonal entries of the confusion matrix indicate that the proportion of wall voxels misclassified as voids rises from 14% to 31% when comparing the best base and CL models. At large intertracer separations, the model predictions on the edges of voids degrade, as seen in the red pixels in Figure 6. Furthermore, very small regions of walls are misclassified as voids, as shown in the blue regions of Figure 6.
@@
PR curves take advantage of the full distribution of class-specific activation scores. For example, to compute the void PR curve, we vary the void class score above which we label voxels as void (note that this assignment differs from the prediction of classes described in Section 4.1 above, where a voxel is assigned the class with the highest score).
@@
For each void classification threshold, we compute the precision – defined as the proportion of correctly identified void voxels out of all predicted void voxels – and the recall – defined as the fraction of true void voxels identified. Varying the class threshold along the PR curve, one obtains perfect precision at extremely high threshold (false positives approach zero), and perfect recall at extremely low threshold (false negatives approach zero). F1 scores (grey curves in Figure 8) measure the balance between precision and recall. Appendix B provides further details of the precision and recall metrics.
[…]
@@
Figure 8: Precision-Recall (PR) curves of void and wall classification for base and CL models with intertracer spacings $\lambda=0.33$ and $\lambda=10$, respectively. Grey lines represent lines of constant F1 score (harmonic mean of precision and recall) for reference. The legend includes the average precision (AP) of each case. The models’ performance on the void class across the large difference of intertracer separation is remarkably consistent. However, performance on wall classification degrades significantly, as seen in this figure and in the confusion matrix (Figure 7).
8 Conclusions

### v2

@@ V.4 Prediction
DeepVoid computes class scores for void, wall, filament, and halo classes that sum to unity in each voxel and assigns the class with the highest score as the prediction for each voxel (see Section IV.3). Due to GPU memory constraints, we cannot load and perform a prediction on an entire TNG mass density cube. Instead, we split the volume into overlapping subcubes of size $128^{3}$ that overlap by half, similar to the training process. DeepVoid then predicts in batches of eight subcubes. To reconstruct the full predicted volume, we reassemble the overlapping subcubes by selecting only the central region of each predicted subcube. This recombination process produces a seamless, nonoverlapping reconstruction of the predicted semantic segmentation with consistent predictions across subcube boundaries.
DeepVoid offers a significant speedup compared to the FFT-based calculation of the tidal tensor segmentation. Prediction of a $512^{3}$ cube takes 15 minutes on a V100 Tesla card versus 88 minutes on a Xenon server with 24 CPUs. This represents an improvement of almost 600% which makes DeepVoid a more efficient alternative to traditional analysis. However, we note that the ultimate purpose of DeepVoid is not to speed up estimation of the tidal tensor segmentation but rather to enable the use of the predicted tidal tensor for deep galaxy redshift surveys with artifacts in which we do not have sufficient sampling density to use standard techniques.
@@
VI Void-Finding Analysis
To quantitatively evaluate the performance of the DeepVoid models, we employ a combination of standard multilabel classification metrics. All of these metrics are calculated from the confusion matrix (as shown below in Figure 7). The confusion matrix is a fundamental tool for evaluating classifiers that compares predicted class labels against the truth labels, so diagonal elements in the matrix represent correct classifications, while the nondiagonal elements represent the instances where the predicted and true class do not match. Definitions and a more detailed discussion of these metrics can be found in Appendix B.
To evaluate model performance, we focus on the void-class F1 score and the Matthews correlation coefficient (MCC) as the primary indicators of model performance. We also consider the balanced accuracy (BA) and the micro-averaged F1 score. The BA and F1 scores range from zero to one, while the MCC ranges from 1-1 to 1, where a zero MCC corresponds to random guessing. These metrics were selected because of their interpretability and robustness to class imbalance.
VI.1 DeepVoid Base Models
As discussed above, we train base models on the full-density TNG-300-3-Dark simulation, with interparticle separation of $\lambda=0.33$ $h^{-1}$ Mpc. The best base TNG model achieves a void F1 score of 0.96 and an MCC of 0.81. Table 1 shows performance metrics of the best DeepVoid base models trained on the TNG DM particles and using SCCE loss functions. The upper row of Figure 5 compares the prediction of the best base model with the truth table from the tidal tensor classification; there is excellent visual agreement. In Figure 5 and throughout this paper, when we plot slices through the prediction cubes, we include voxels that are in both validation and training data so that one can see the contiguous structures. However, only validation data are used for the calculation of performance metrics. In Section VII.3 below, we quantify details of this visual comparison in the context of examining the effects of sparse sampling on the model performance.
@@ VI.1 DeepVoid Base Models
VI.2 Alternative Validation Test
As an additional test, we rotate the DM particles/halo distribution in TNG300 by 45 and recompute the tidal tensor mask. CNNs are by nature not rotationally invariant (Zeiler and Fergus, 2013).
@@ VI.2 Alternative Validation Test
VI.3 Toward Void Catalogs
At this stage of the development of DeepVoid we have chosen not to produce “void catalogs” that divide the void regions into a list of discrete voids. The tidal tensor method for segmenting large-scale structure into voids, walls, filaments, and halos does not by itself provide a “truth table” of discrete structures for training the network. For voids, we know that the interior structure of voids is complex: thus the identification of individual voids requires one to specify a spatial scale on which to segment these inherently multiscale structures (Aragon-Calvo and Szalay, 2013; Aragon-Calvo, 2024). Geometric void finders such as, for example, VIDE (Sutter et al., 2015), REVOLVER (Nadathur and Hotchkiss, 2014), and VoidFinder (El-Ad and Piran, 1997; Hoyle and Vogeley, 2002) implicitly or explicitly choose a scale that is determined by the average intergalaxy spacing; below some scale, the observed underdensities are not statistically significant compared with Poisson fluctuations in the galaxy density, and so these methods ignore small-scale voids by default or by choice. In future work, we will examine criteria for creating discrete void catalogs from DeepVoid void regions.
@@ VI.3 Toward Void Catalogs
Figure 5: Upper row: TNG mass density contrast and prediction for the best base TNG model predicting on the full TNG particle distribution with $\lambda=0.33$ $h^{-1}$ Mpc. Upper right panel is the tidal tensor truth table, which is the same for all models. Lower row: TNG subhalo positions and prediction for the best CL model. Subhalos are first ordered by mass, then abundance matched such that the average intertracer separation is $\lambda=10$ $h^{-1}$ Mpc. Subhalos that are within 2 $h^{-1}$ Mpc of either side of the slice taken through the simulation are shown. The base model was trained on the TNG full density with a U-Net with three layers of depth and 32 initial filters. The model was then curriculum-trained on subhalos with $\lambda=3$ $h^{-1}$ Mpc average spacing and then curriculum-trained again out to $\lambda=10$ $h^{-1}$ Mpc. While the slices through the prediction cubes necessarily include voxels in both the training and validation data, only the validation data are used to compute performance metrics reported in this paper.
VII Training on Sparser Samples
The DeepVoid base model performs quite well, indicating that training has produced sets of convolutional filters that identify the spatial pattern of structures that result from gravitational physics predicted by the tidal tensor. In that first step of the DeepVoid approach, the performance of the base model indicates that the DeepVoid model is an accurate “detector” of the predicted shapes of voids and other structures in the full matter density field, for which DeepVoid acts as an emulator of the tidal tensor segmentation. For application to observational data, the next step is to extend the model to detect the pattern of tidal tensor structures in the distribution of galaxies, which are sparse, biased tracers of the matter density.
@@ VII Training on Sparser Samples
Thus, to explore building a version of DeepVoid that can be applied to density fields from deep galaxy surveys, we must train and predict on density fields with larger intertracer spacings than the full DM particle fields of TNG. The subhalo catalog of TNG provides an adequate test case for DeepVoid, as described in Section II.
VII.1 Directly Training with Sparse Samples
@@
The most naive way to predict the morphology of a volume with a large average intertracer separation would be to simply run a prediction using a base model. As expected, the base models, which are trained on density fields with intertracer spacings of much less than one, perform very poorly when predicting on the $\lambda=10$ $h^{-1}$ Mpc subhalo mass density. The best TNG base model reaches a BA of 0.27 and an MCC of 0.07. Recalling the definition of the MCC, that score is only slightly better than random guessing.
@@
Next, we train a new model directly on the volumes with intertracer separations similar to those of galaxy redshift surveys of interest. This vastly outperforms the base models. Training on the subhalo density field with $\lambda=10$ $h^{-1}$ Mpc results in scores of void F1=0.89 and MCC=0.56. As discussed in the next section, we can further improve the models’ predictive performance using a multistep training process.
@@
Figure 6: Slices through the TNG300-3-Dark volume. The tidal tensor mask is shown in gray-scale on both panels. Red voxels represent true void voxels misclassified as wall, while blue voxels show true wall voxels misclassified as void. Left: base model trained on TNG dark matter particles with a 0.33 $h^{-1}$ Mpc separation. Right: CL model trained on subhalos with a 10 $h^{-1}$ Mpc spacing. As the intertracer separation increases, the lower signal-to-noise ratio negatively impacts the quality of the segmentation, causing the model to miss small-scale features such as walls between smaller voids and tendrils of filaments extending into voids. Note that while it may appear that the percentage of red voxels does not significantly increase in the right panel, as we would expect to see from the confusion matrix in Fig. 7 where 4% to 12% are reported for the left and right panels, this appearance is due to the particular slice shown. Also, while these slices through the prediction cubes necessarily include voxels in both the training and validation data, only the validation data are used to compute performance metrics reported in this paper.
@@
Figure 7: Comparison of two confusion matrices generated by running model predictions on the TNG validation dataset. Left: confusion matrix from a model trained with depth of 3 and 32 initial filters on the full DM density (average interparticle spacing of 0.33 $h^{-1}$ Mpc). Right: confusion matrix from a model with depth of 3 and 32 initial filters predicting on TNG subhalos with an average intertracer spacing of 10 $h^{-1}$ Mpc. This model was originally trained on the full TNG DM density, then curriculum-trained out to 3 $h^{-1}$ Mpc, and again to 10 $h^{-1}$ Mpc. The numbers in the matrices are normalized by the true instances, such that the rows of each matrix sum to one. The base model achieves a void F1 score of 0.96 and a MCC of 0.81 while the CL model achieves a void F1 score of 0.89 and a MCC of 0.6.
VII.2 Applying Curriculum Learning
@@
Note. — DeepVoid models that used curriculum learning to extend out to an intertracer separation of 10 $h^{-1}$ Mpc. As always, all listed scores were generated from predicting on the validation data. The type of curricular learning employed in each training is denoted by the column CL Type: ENC_EO meaning every other convolution on the entire encoding side is frozen, ENC_D1 and ENC_D2 representing all convolutions on the encoder side being frozen down to a depth of two and three, respectively. A CL type of ENC_D1_ENC_D2 means that the model has been curricular trained twice, once to an intermediate intertracer spacing (here, $\lambda_{int}=$ 3 $h^{-1}$ Mpc) then again to 10 $h^{-1}$ Mpc. $\lambda_{\rm base}$ represents the interparticle or intertracer spacing of the density field that the base model was trained on, while $\lambda_{\rm int}$ is the intermediate intertracer spacing, if there is one for each respective curricular learning.
As outlined in Section V.3, curriculum learning aims to successively train a model on increasingly difficult examples. Typically, when training a network, the weights of the model are initialized randomly and then refined iteratively through the process of stochastic gradient descent. However, this is not the case for curriculum learning, as some of the initial weights of the model have been preserved from previously trained models that were trained on a simpler example of the task. The idea is that the model inherits valuable knowledge from earlier stages of training, then further iterates to an even better region of the loss landscape. In practice, this means that when we introduce a more difficult example to the model, such as a sparser mass density field, we freeze some of the weights that were learned by training on the ‘easier’ (in this case, denser) example.
We experiment with different schemes for freezing the weights. Commonly, the weights of the encoding side of the U-net are frozen to allow for a stable feature extraction process, while the decoding side is allowed to focus on interpreting the already learned features and applying them to the new, more challenging data (Ardalan and Subbian, 2022). We perform curriculum learning using three different schemes: freezing weights in the entire encoding side (ENC), freezing weights in only one of the two convolutional blocks on the encoding side (ENC_EO), or freezing weights only down to some layer of depth on the encoding side (ENC_D1/D2). For each CL scheme, we begin with weights from the best base model, freeze the specified set of weights, and train the unfrozen weights using training data with an intertracer separation that is larger than that on which the base model was trained. No models with the ENC scheme outperformed the ENC_EO models, so they are not discussed further in this analysis.
@@
For every model in this table, $\lambda_{\rm pred}=10$ $h^{-1}$ Mpc. The best CL models achieve a void F1 score of 0.89 and an MCC of 0.60. The two-step model slightly outperforms one-step CL models in BA. The lower row of Figure 5 shows the result of training a model using this two-step CL scheme out to intertracer separation of 10 $h^{-1}$ Mpc. As with the base model, there is excellent visual agreement between the predicted class labels and the ground truth derived from the tidal tensor. In Section VII.3 we quantify details of this comparison.
Note that Tables 1 and 2 present only the most successful training schemes. We experimented with numerous other combinations of hyperparameters and loss functions, not shown here. These results represent a successful proof-of-concept test of classifying LSS from a sparse distribution, but are by no means exhaustive.
Model BA Micro F1 Void F1 MCC
Base Binary 0.94 0.95 0.96 0.88
Base Multiclass 0.78 0.91 0.96 0.81
CL Binary 0.81 0.85 0.89 0.66
CL Multiclass 0.70 0.81 0.89 0.60
Table 3: Comparison of classification scores: Balanced Accuracy (BA), Micro-averaged F1, Void class F1 and Matthews correlation coefficient (MCC) for Base and Curriculum Learning (CL) UNet models. CL scores are for $\lambda=10$ Mpc/h/h.
From these results, we conclude that performing CL improves the quality of the segmentation for sparser samples. In the case of voids (see Void F1 score) both binary and multiclass classifications perform identically, even at high subsampling (see Table 3 for a comparison between scores for base and $\lambda=10$ models with and without CL). MCC and Micro F1 scores are slightly lower for multiclass models because small misclassifications within the nonvoid classes (e.g., filament vs. wall) reduce those metrics. Using CL techniques, we improve training on subhalo distributions with intertracer spacings as large as $\lambda_{\rm pred}=10$ $h^{-1}$ Mpc. We find that the two-step CL scheme (ENC_D1_ENC_D2) yields slightly better results than a single-step CL scheme. Further exploration of model hyperparameters might yield further improvements.
VII.3 Segmentation Performance in Sparse Samples
@@
An increase in the number of misclassified voxels with a reduction in tracer density is highlighted in the slice plots of Figure 6. Voxel misclassification is quantified in the confusion matrix in Figure 7, where the left panel describes results for the base model trained on the full DM distribution in TNG300 and the right panel describes results for CL training on very sparse subhalos. The off-diagonal entries of the confusion matrix indicate that the proportion of wall voxels misclassified as voids rises from 14% to 31% when comparing the best base and CL models. At large intertracer separations, the model predictions on the edges of voids degrade, as seen in the red pixels in Figure 6. Furthermore, very small regions of walls are misclassified as voids, as shown in the blue regions of Figure 6. Note that the exact percentages of misclassification in the particular slice shown in Figure 6 do not match the statistics over the entire simulation volume, simply due to variation in the small number of such structures in a single thin slice.
@@
PR curves take advantage of the full distribution of class-specific activation scores. For example, to compute the void PR curve, we vary the void-class score from 0 to 1, using the same score across the entire simulation cube to separate void voxels from nonvoid voxels. This separation of void voxels from nonvoid voxels is different from the multiclass segmentation above where each voxel was assigned the class with the highest score for that voxel alone.
@@
For each void-class score used to separate void voxels from nonvoid voxels, we compute the precision – defined as the proportion of correctly identified void voxels out of all predicted void voxels – and the recall – defined as the fraction of true void voxels identified. Varying the class threshold along the PR curve, one obtains perfect precision at an extremely high threshold (false positives approach zero), and perfect recall at an extremely low threshold (false negatives approach zero). F1 scores (grey curves in Figure 8) measure the balance between precision and recall. Appendix B provides further details of the precision and recall metrics.
The CL model accurately classifies void voxels even with a large increase in intertracer spacing. This is shown by the relatively stable void PR curve, which indicates that the model remains effective at void identification, with the CL-trained model achieving an average precision $\text{AP}=0.91$ and $\text{F1}=0.89$ (harmonic mean of precision and recall) for voids in sparse data ($\lambda=10$ $h^{-1}$ Mpc). In contrast, the average precision for walls drops significantly compared to the base model. The corresponding confusion matrices suggest that this drop in wall precision reflects an increase in the fraction of voids predicted as wall from 4% to 12%, and a rise in the fraction of filaments predicted as wall from 22% to 36%. Thus, the PR curves and confusion matrices quantify the degradation of predictions along void boundaries observed in Figure 6.
@@
Figure 8: Precision-Recall (PR) curves of void and wall classification for base and CL models with intertracer spacings $\lambda=0.33$ $h^{-1}$ Mpc and $\lambda=10$ $h^{-1}$ Mpc, respectively. Grey lines represent lines of constant F1 score (harmonic mean of precision and recall) for reference. The legend includes the average precision (AP) of each case. The models’ performance on the void class across the large difference of intertracer separation is remarkably consistent. However, performance on wall classification degrades significantly, as seen in this figure and in the confusion matrix (Figure 7).
Figure 9: Density distribution of voxels inside voids, walls, filaments, and clusters. The training mask is shown in grey while the predictions for the $\lambda=3$$h^{-1}$ Mpc and $\lambda=10$$h^{-1}$ Mpc samples are indicated by red and blue lines respectively.
Figure 9 shows a comparison between the density distribution inside the true and predicted masks. It serves as a first test of the physical validity of our results. The density distributions follow the expected, behavior increasing from voids to walls, filaments and clusters. While each cosmic environment has a distinct peak, there is a large overlap between the distributions that highlights the difficulty in separating environments by density alone. The stable performance of DeepVoid, shown in the similarity of the three samples, even in the sparsest sample, is encouraging. However, as shown in Figs. 6 and 7, at low sampling density ($\lambda=10$ $h^{-1}$ Mpc), some void voxels near the peak and at slightly lower density are misclassified as relatively low-density wall voxels, while some wall voxels at density above the wall peak are misclassified as relatively high-density void voxels. This is not unexpected, as the low sampling decimates the structures DeepVoid is seeking to identify.
VIII Conclusions